# Unreal Robotics Lab: A High-Fidelity Robotics Simulator with Advanced Physics and Rendering

### v2

@@ III-G Example Deployment of Robots, Controllers, and Visual Planners
III-G1 Example Robot Integrations
@@ IV-A Simulation and Real-World Comparision
To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM, EigenCAM, cosine similarity and KL Divergence. These methods assess both spatial attention and feature distribution similarity. Grad-CAM computes the gradient of the class score $y^{c}$ with respect to feature map activations:
@@
where $\mu$ is the mean activation and $v_{1}$ is the first principal component.
@@ V CONCLUSION
This paper presents a high-fidelity robotics simulation framework that integrates advanced photorealistic rendering with precise physics modeling to support vision-based robotics research and benchmarking. Using Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions. Experimental results confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios. Despite its strengths, the system has certain limitations, including the absence of direct reinforcement learning (RL) support, lack of deformable terrain simulation, limited scene generation capabilities, and a focus on general robotics rather than autonomous vehicles or high-precision drone simulations. To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [zakka2025mujocoplayground]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation. Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.

### v3

@@ III-F Example Deployment of Robots, Controllers, and Visual Planners
III-F1 Example Robot Integrations
@@ IV-A Simulation and Real-World Comparision
To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM [selvaraju2017grad], EigenCAM [muhammad2020eigen], cosine similarity and KL Divergence [kullback1951information]. These methods assess both spatial attention and feature distribution similarity. Specifically, spatial attention is computed using Grad-CAM and EigenCAM in equations 1 and 2 respectively to assess feature distribution similarity. Grad-CAM (Gradient-weighted Class Activation Mapping) computes gradients of class score $y^{c}$ (pre-softmax logit for class $c$) w.r.t. feature maps $A^{k}$ (spatial activations from channel $k$ in the final conv layer):
@@
where $\alpha_{k}^{c}$ is the Global Average Pooling (GAP)-weighted importance of channel $k$ for class $c$ ($i,j$: spatial indices; $Z$: width $\times$ height). Heatmap $L^{c}$ is the weighted sum of maps with ReLU, yielding a saliency map of regions influencing the model’s class $c$ prediction.
@@
where $A_{ij}$ is the activation vector across channels at position $i,j$; $\mu$ is its mean over spatial locations, and $v_{1}$ is the first principal component (eigenvector of $C$). The map $L$ is the projection onto $v_{1}$.
@@ V CONCLUSION
This paper presents a high-fidelity robotics simulation framework that integrates advanced photorealistic rendering with precise physics modeling to support vision-based robotics research and benchmarking. Using Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions. Experimental results, such as utilizing the Replay System to test varied visual effects on identical trajectories, and Niagara-driven adversity benchmarks highlighting planner failures in volumetric smoke, confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios.
Despite its strengths, the system has certain limitations, including the lack of deformable terrain simulation, one-way kinematic coupling for UE driven Actors, and the absence of validated sensor noise models for depth and LiDAR data. To address these limitations and further improve our simulator, we plan to focus on narrowing the sim-to-real perceptual gap by implementing procedural material randomization, generate large-scale synthetic datasets to enhance vision-based model training, introduce VR teleop for data generation, expand automated scene generation, and enable bilateral physical interactions between MuJoCo and UE. Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.